# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### v2

@@ 5.1 Experimental Setup
  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects. The retrieved objects are reranked by ${}_{D}$. We set the number of clusters $c$ to be the same as our index. Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4. Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
Table 3: Comparison of retrievers across three datasets.
@@
Recall NDCG Recall NDCG Recall NDCG
@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
TkQ+${}_{D}$
0.7263 0.6633 0.5056 0.4145 0.7717 0.7348 0.5742 0.4556 0.7357 0.7012 0.5797 0.4878
IVF+${}_{D}$
0.5461
0.4954
0.3771
0.3103
0.4965
0.4717
0.3713
0.3000
0.4720
0.4535
0.3835
0.3274
[…]
0.4126
0.3714
0.3025
0.2415
0.4174
0.3812
0.2960
0.2101
0.3921
0.3714
0.3011
0.2419
LSH+${}_{D}$
0.5473
0.5057
0.3973
0.3385
0.5356
0.5097
0.4135
0.3453
0.6686
0.6433
0.5445
0.4661
HNSW+${}_{D}$
0.6267
0.5828
0.4545
0.3765
0.5527
0.5270
0.4182
0.3432
0.5009
0.4807
0.4066
0.3468
IVFPQ+${}_{D}$
0.7064
0.6400
0.4872
0.4002
0.7184
0.6765
0.5257
0.4188
0.6997
0.6724
0.5650
0.4791
LIST 0.7711 0.7170 0.5668 0.4812 0.7721 0.7401 0.6099 0.5139 0.8250 0.7909 0.6747 0.5815
5.63%
7.95%
10.70%
15.32%
0.05%
0.72%
6.21%
12.54%
12.13%
12.79%
16.38%
19.21%
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx. In our model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 . The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets. $\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section 5.3. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
5.2 Effectiveness Study (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table 2 reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW and ${}_{D}$. This can be attributed to the word mismatch issue discussed in Section 3.2. Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by liu2023effectiveness . (3) Directly applying the dual-encoder model to our task is less effective. On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance. DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions DBLP:conf/sigir/DingCXHLZX23 .
[…]
Table 4: Query runtime on three datasets (ms).
Beijing
Shanghai
Geo-Glue
TkQ
6.3
6.5
1,178.0
TkQ+${}_{D}$
86.3
84.9
5,412.3
IVF+${}_{D}$
2.9 2.9 4.0
LSH+${}_{D}$
5.2
5.1
31.1
HNSW+${}_{D}$
7.3
6.5
14.6
IVFPQ+${}_{D}$
3.6
3.7
5.2
LIST 3.0 3.0 5.1
5.3 Efficiency Study (RQ2)
Runtime of LIST. We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines. During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory. The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set. We present the total runtime of different methods in Table 4. Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST. However, their effectiveness is much worse than LIST as stated in Table 3. Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets. Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets.
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Figure 4: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters ($c$) (up and right is better).
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters ($c$). We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings. Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100. For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter the number of objects retrieved (top-$k$) from 40,000, 20,000, 1,$0$, 5,000 to 2,000. This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison. Other hyperparameter settings are fixed. We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results. We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency. This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed. (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly. This is due to the word mismatch issue discussed in Section 3.2. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+${}_{D}$ exhibits almost constant results. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$). We further investigate the impact of the number of clusters $cr$ to route queries and objects. We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$. For other baselines, we set hyperparameter the number of objects retrieved (top-$k$) from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study. Other hyperparameter settings are fixed. Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5. According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings. Increasing $cr$ can enhance effectiveness, but it also leads to higher latency. In practice, users can adjust this parameter according to their needs.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table 5, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
Table 5: Memory usage on three datasets (MB).
Beijing
Shanghai
Geo-Glue
TkQ+${}_{D}$
719
857
12,037
IVF+${}_{D}$
508
505
8,510
LSH+${}_{D}$
513
511
8,638
HNSW+${}_{D}$
548
545
9,427
IVFPQ+${}_{D}$
508
505
8,520
LIST
508
505
8,515
5.4 Scalibility Study (RQ3)
Scalability Study. We further explore the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. Since there is no ground-truth data for the relevance between the newly crawled POIs and the existing queries, we only report the efficiency of our proposed retriever LIST and our relevance model LIST-R. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 6. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.
(a) Scalability Study of LIST-R
(b) Scalability Study of LIST
Figure 6: Scalability study on Geo-Glue dataset.
@@ 5.5 Ablation Studies (RQ4)
Index Study. To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset. We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects. The results are reported in Table 6, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings. We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table 4), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1. (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed. The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table 4.
Table 6: Index ablation study on the Geo-Glue dataset.
Recall NDCG
@20 @10 @5 @1
TkQ+LIST-R
0.7890
0.7585
0.6517
0.5639
IVF+LIST-R
0.5381
0.5192
0.4671
0.3960
LSH+LIST-R
0.7665
0.7381
0.6378
0.5555
HNSW+LIST-R
0.7527
0.7231
0.6219
0.5390
LIST 0.8250 0.7909 0.6747 0.5815
@@
LIST-R + $S_{in}$
0.7526
0.7087
0.5255
0.4271
LIST-R + $\alpha*S_{in}^{\beta}$
0.5308
0.4532
0.3130
0.2411
LIST-R + ADrW
0.7925
0.7414
0.5832
0.4792
LIST-R
0.8156 0.7545 0.5913 0.4989
@@
Effect of Training Dataset Size on Effectiveness. We investigate the impact of training dataset size over LIST and LIST-R. Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process. During testing, we use the complete test dataset. This approach avoids the issue of data leakage. We vary the percentage of objects from 30%, 50%, 80% to 100%. We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results. The results are shown in Figure 8(a). LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval. The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
(a) The Impact over Effectiveness
(b) The Impact over Training time
Figure 8: The impact of training dataset size over effectiveness and training time on the Beijing dataset.
In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
Effect of Training Dataset Size on Training Time. We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 8(b). Other datasets exhibit similar trends. We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.

### v3

@@ 5.1 Experimental Setup
  • IVFPQ+LIST-R: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects. The retrieved objects are reranked by LIST-R. We set the number of clusters $c$ to be the same as our index. Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 2. Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by LIST-R. To ensure a fair comparison, we set $k$ to 5,000 for the baselines that involve $k$ on the Beijing and Shanghai datasets, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 30,000 for the same reason above. To investigate the capability of LIST to trade off effectiveness and efficiency, we evaluate the effect of the hyperparameter $cr$ and $k$ in Section 5.3.
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx. In our proposed relevance model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 . The hyperparameter settings of LIST are detailed in Table 2. Specifically, the spatial footstep controller $t$ of $\mathcal{T}$ is set to 1000 for all three datasets. The hyperparameters to control the generation of pseudo labels, $\text{neg}_{\text{start}}$ and $\text{neg}_{\text{end}}$, are set to 50,000 and 55,000 for the Beijing dataset, 60,000 and 65,000 for the Shanghai dataset, and 180,000 and 181,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ settings shown in Section 5.3. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: V100 32GB.
(a) Beijing
(b) Shanghai
(c) Geo-Glue
Figure 4: The effectiveness-efficiency trade-off results (upper and right is better).
5.2 Effectiveness of Proposed Relevance Model (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table 3 reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 31.60% improvement over the best baseline on NDCG@1 and up to a 59.92% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW. This can be attributed to the word mismatch issue discussed in Section 3.2. Notably, PALM is outperformed by TkQ, which is consistent with the results reported by liu2023effectiveness .
(a) Beijing
(b) Shanghai
5.3 Effectiveness-Efficiency Trade-off of Proposed Index (RQ2)
Effectiveness-efficiency trade-off of LIST. We investigate the effectiveness-efficiency trade-off of LIST by comparing it with the state-of-the-art index baselines, which all use our relevance model. The evaluation results are shown in Figure 4, where LIST-R represents the brute-force search using our relevance model. We have the following findings: (1) Overall, LIST consistently outperforms all baselines, offering the best trade-off between effectiveness and efficiency across all three datasets. Specifically, compared to LIST-R, LIST is an order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. While most ANNS indexes have similar runtime to LIST, their effectiveness falls short. (2) Directly applying ANNS indexes to spatial keyword queries results in a significant drop in effectiveness. Compared to LIST-R, which uses brute-force search, existing ANNS indexes sacrifice considerable effectiveness for improved efficiency. This could be attributed to the fact that these ANNS indexes do not consider the spatial factor. (3) When TkQ is used as an index baseline, it is significantly slower. The reason is that the IR-Tree is not designed to retrieve a large number of objects. As the hyperparameter $k$ increases, the IR-Tree’s pruning ability declines, which is consistent with previous empirical results congEfficientRetrievalTopk2009 . (4) Simply modifying the existing ANNS index to incorporate the spatial factor is even less effective. IVF+LIST-R consistently outperforms IVFS+LIST-R on the three datasets. This confirms that manually assigning a weight to balance the two factors is ineffective.
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$). We examine LIST’s ability to trade off effectiveness and efficiency. Here, we vary the hyperparameter $cr$, which represents the number of clusters to route, from 1, 2, to 3 on the Beijing and Shanghai datasets. For baselines that share the same hyperparameter with LIST, i.e., IVF+LIST-R, IVFS+LIST-R, and IVFPQ+LIST-R, we maintain the same increment for $cr$. For other methods, we adjust the hyperparameter $k$, which represents the number of objects retrieved by the indexes, to be close to the number of objects retrieved by our proposed index, adjusting it from 5,000, 10,000, to 20,000 on the Beijing and Shanghai datasets. For TkQ+LIST-R, due to its slow query processing speed, we adjust $k$ from 100, 500, 1,000, 5,000, and 10,000 to 20,000. We present the trade-off results in Figure 5. We have the following findings: (1) LIST consistently outperforms other methods, providing a better trade-off between effectiveness and efficiency. (2) Compared to other methods, the effectiveness of TkQ+LIST-R increases more slowly as $k$ increases, while the time overhead rises significantly. This is due to the word mismatch issue discussed in Section 3.2. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+LIST-R is almost not affected by the number of retrieved objects. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table 4, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+LIST-R, TkQ+LIST-R, and HNSW+LIST-R, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
Table 4: Memory usage on three datasets (MB).
@@ 5.3 Effectiveness-Efficiency Trade-off of Proposed Index (RQ2)
TkQ+LIST-R 719 857 12,037
IVF+LIST-R 508 505 8,510
LSH+LIST-R 513 511 8,638
HNSW+LIST-R 548 545 9,427
IVFPQ+LIST-R 508 505 8,520
LIST 508 505 8,515
(a) Beijing
(b) Beijing
Figure 6: The effectiveness-efficiency trade-off results for different relevance models (upper and right is better).
5.4 Generalization Study of the Proposed Index (RQ3)
Generalization Study. We explore the applicability of our proposed index to other relevance models designed for embedding based spatial keyword queries. Specifically, we introduce two variants of our relevance model: LIST-RRoberta and LIST-RLinear. Specifically, in LIST-RRoberta, the BERT model in the dual-encoder module is replaced by another pre-trained language model RoBERTa [^5] https://huggingface.co/clue/roberta_chinese_base wolf2019huggingface . In LIST-RLinear, the learned spatial relevance module is replaced with the linear function of distance. We train the two baselines on the Beijing dataset and use them to generate embeddings in advance. We then use our proposed index to build an index on these embeddings and compare it with existing index baselines. The comparison of our proposed index with the baseline indexes is shown in Figure 6. The experimental results demonstrate that our proposed index maintains a similar advantage over baseline indexes when applied to other relevance models.
5.5 Scalibility Study (RQ4)
Scalability Study. We evaluate the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. Since there is no ground-truth data for the relevance between the newly crawled POIs and the existing queries, we only report the efficiency of our proposed retriever LIST and our relevance model LIST-R. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 7. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.
(a) LIST-R
(b) LIST
Figure 7: Scalability study on the Geo-Glue dataset.
@@ 5.6 Ablation Studies (RQ4)
LIST-R + $S_{in}$ 0.7526 0.7087 0.5255 0.4271
LIST-R + $\alpha*S_{in}^{\beta}$ 0.5308 0.4532 0.3130 0.2411
LIST-R + ADrW 0.7925 0.7414 0.5832 0.4792
LIST-R 0.8156 0.7545 0.5913 0.4989
@@
Effect of Training Dataset Size on Effectiveness. We investigate the impact of training dataset size over LIST and LIST-R. Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process. During testing, we use the complete test dataset. This approach avoids the issue of data leakage. We vary the percentage of objects from 30%, 50%, 80% to 100%. We present the results for NDCG@1 and Recall@10 on the Geo-Glue dataset using default hyperparameters, and similar results are observed for the remaining metrics and datasets. The results are shown in Figure 10. LIST-R uses brute-force search with the relevance model and LIST utilizes our proposed index for retrieval. The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness. Additionally, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
Effect of Training Dataset Size on Training Time. We also measure the training time per epoch of LIST-R and LIST on the Geo-Glue dataset by varying the size of the training data in the same manner stated in the last paragraph, and the results are shown in Figure 9. Other datasets exhibit similar trends. We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
Figure 9: The impact of training dataset size on training time for the Geo-Glue dataset.
Figure 10: The impact of training dataset size on effectiveness on the Geo-Glue dataset.
Effect of Training Sample Selection on Effectiveness. To investigate the impact of training sample selection on effectiveness, we select different 50% subsets from the Geo-Glue training dataset for training and then compare their effectiveness to demonstrate the influence of training sample choice. Since the random selection is determined by a random seed, we use different seeds (e.g., 1, 2, 3, 4) to ensure that each selection of training samples is different. The experimental results on the Geo-Glue dataset are presented in Figure 11, where we only report the results for NDCG@1, and the other metrics are similar. The results show that LIST and LIST-R produce similar results across different training samples, demonstrating that our solution is robust to variations in training sample selection.
Figure 11: The impact of training sample selection on effectiveness for the Geo-Glue dataset.
Figure 12: Results of Low-Frequency Descriptors.
Evaluating Queries Containing Keywords of Low Frequency. To investigate the capability of LIST and LIST-R in handling unique or low-frequency keywords, such as ’halal’ or ’vegan’, which are more distinctive than common keywords like ’pizza’ or ’pasta’, we investigate whether the frequency of query keywords affects the performance of LIST and LIST-R on the geo-glue dataset. Specifically, we calculate the frequency of each keyword and identify queries containing low-frequency keywords, along with their corresponding ground-truth labels. The frequency of query keywords in the geo-glue dataset ranges from 1 to 3,891. We select queries containing keywords that appear at most $freq$ times, with $freq$ set to 1, 10, 100, and 1,000, resulting in subsets of test queries consisting of 4,108, 15,892, 19,930, and 19,991 queries, respectively. We then evaluate LIST and LIST-R on these subsets to demonstrate their capability in handling queries containing low-frequency keywords. The experimental results on the geo-glue dataset are shown in Figure 12, where we only report the results for Recall@10 and NDCG@1, as similar trends are observed on other metrics. The results show that LIST and LIST-R maintain performance for queries containing keywords of low-frequency.